OpenStudy (anonymous):

Simplify(2sqrt(6)+7sqrt(5))(2sqrt(10)-5sqrt(3))

OpenStudy (anonymous):

\[2\sqrt{6}+7\sqrt{5} *2\sqrt{10}-5\sqrt{3}\] right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Is that the simplest form

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\sqrt{a}*\sqrt{b}=\sqrt{ab}\] apply it

OpenStudy (anonymous):

Would the whole thing be in sqrt

OpenStudy (anonymous):

use it for the second term and then you will use it to the others

OpenStudy (anonymous):

\[=2\sqrt{3*2}+14\sqrt{50}-5\sqrt{3}\]

OpenStudy (anonymous):

Could you show how to do it step by step

OpenStudy (anonymous):

then \[2\sqrt{3}\sqrt{2}+14\sqrt{25*2}-5\sqrt{3}\]

OpenStudy (anonymous):

\[2\sqrt{3}\sqrt{2}+14\sqrt{25}\sqrt{2}-5\sqrt{3}\]

OpenStudy (anonymous):

Is that the finale answer

OpenStudy (anonymous):

\[2\sqrt{3}\sqrt{2}+14*5\sqrt{2}-5\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{2}(2\sqrt{3}+70)-5\sqrt{3}\]

OpenStudy (anonymous):

Final answer

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

anytime :)